OpenStudy (anonymous):

what are the solutions to 2y^2-16=8y

OpenStudy (anonymous):

not factorable?

OpenStudy (anonymous):

i dont think so

OpenStudy (anonymous):

it would not come out even and integer, but the solutions are 2(1-sqrt(3)), and 2(1+sqrt(3))

OpenStudy (anonymous):

2y2-8y-16=0 y2-4y-8=0 y= 4+- sqrt(16+32) /2 y= 4+- 4 sqrt3 /2 y= 2+- 2 sqrt3

OpenStudy (anonymous):

solve b using the quadradic equation

OpenStudy (anonymous):

Ya that's what I did

OpenStudy (anonymous):

thank you!